# Bidirectional Decoding: Improving Action Chunking via Closed-Loop Resampling

### v1

@@ 4 Method: Bidirectional Decoding
As analyzed in Sec. 3, action chunking facilitates the modeling of temporal dependencies in demonstrations but sacrifices reactivity to unexpected states in stochastic environments. In this section, we address this issue by bridging long action chunks with closed-loop operations. We will first outline the general framework in Sec. 4.1 and then describe two specific criteria in Sec. 4.2.
@@ 4.1 Test-Time Search
Given a generative policy with context length $c$ and prediction horizon $l$, an action chunk sampled at time $t$
$a\sim\pi_{\theta}(a_{t},a_{t+1},\cdots,a_{t+l}|{\leavevmode\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{s_{t-c},s_{t-c+1},\cdots,}}s_{t})$ (5)
is expected to adhere to a consistent latent strategy over the next $l$ time steps. However, naive closed-loop operation of the policy entails executing only the first action of each predicted chunk, leading to a sequence of actions $a_{t}^{(t)},a_{t+1}^{(t+1)},\cdots,a_{t+l}^{(t+l)}$ sampled at different time steps. When multiple latent strategies exist in the demonstrations and are learned by the policy (e.g., left versus right, stop versus go, fast versus slow), independently sampled action chunks may oscillate between different strategies, leading to inconsistent behavior that diverges from the demonstrations.
Our main hypothesis is that while the probability of any pair of samples sharing the same latent strategy is low, the likelihood of finding a consistent pair from a large number of samples is significantly higher. This intuition motivates us to cast the problem of closed-loop action chunking as searching for the optimal action among a batch of plans sampled at each time step,
@@
where $\mathcal{A}$ is the set of sampled action chunks, $\mathcal{L}_{B}$ and $\mathcal{L}_{F}$ are two criteria measuring the temporal dependency with respect to the backward decision and forward plan, which we will describe next.
@@ 4.2 Bidirectional Criteria
Backward coherence. To preserve sufficient temporal dependency in closed-loop operations, a sequence of actions should (i) commit to one action chunk in the absence of unexpected changes and (ii) react smoothly to environmental changes. We use the action chunk selected at the previous time step as a reference for enforcing coherence across time. Given the previous action chunk $\hat{a}:={a}^{(t-1)}_{t-1},\cdots,{a}^{(t-1)}_{t+l-1}$, we select action chunks that minimize the weighted sum of Euclidean distances across the $l-1$ overlapping steps:
@@
Here, $\rho$ is a decay hyperparameter to account for growing uncertainty over time. This backward loss encourages similar latent strategies between neighboring steps, while allowing for gradual adaptation to unforeseen transition dynamics.
Forward contrast. An ideal policy should predict far enough into the future to capture the planning capabilities inherent in human demonstrations. However, building such a policy can be challenging in practice due to modeling constraints and dataset limitations. Often, even the best policy available may still produce a significant number of suboptimal plans. To address this, we introduce a forward contrast objective to identify and reject these suboptimal plans. Specifically, we compare each candidate plan with two sets of reference samples: one set from a stronger policy and the other set from a weaker one. We use a well-trained model as the stronger policy, and a model from an early underfitting checkpoint or with a shorter prediction horizon as the weaker policy. Intuitively, the weaker policy cannot capture long-term planning as effectively as the stronger one. Our forward contrast loss is thus framed as minimizing the average distance between a candidate plan and a set of positive samples while maximizing its average distance from the negative ones,
Fig. 4 illustrates the combined effects of the backward coherence and forward contrast criteria on sample selection. Notably, not all samples in $\mathcal{A}^{+}$ and $\mathcal{A}^{-}$ are necessarily subject to the same mode. To mitigate this, we trim each set by removing samples that deviate significantly from the mode of the previous decision. This is achieved by summing over the $K$ smallest distance values for in the positive and negative sets in Eq. 8. The full process of our decoding method is outlined in Algorithm 1. Since all steps in BID can be computed in parallel, the overall computational overhead remains modest on modern GPU devices.
Figure 4: Illustration of bidirectional criteria.
@@ 4.3 Discussions
Interpretation of our method. Our method makes no changes to the learned policy; instead, it intervenes in the model distribution through sample selection. As illustrated in Fig. 11, randomly sampled sequences may be misaligned with both the previous decisions and the target demonstrations. Given a set of candidates, the backward step first identifies the behavioral mode from the past decision stored in memory; the forward step then removes the samples with low likelihood under the target distribution using prior knowledge of positive and negative samples. By comparing samples across time steps and model horizons, our method bridges the gap between the proposal and target distributions during inference.
Relation to recent methods. Our method builds upon the receding horizon [4] and temporal ensembling [3] used in previous works, but with crucial distinctions. Receding horizon seeks a compromise between temporal dependency and dynamic uncertainty by using a moderate action horizon (e.g., half of the prediction horizon), which is inevitably sup-optimal when both factors are prominent. Temporal ensembling strengthens dependency across chunks by averaging multiple decisions over time; however, weighted-averaging operations can be detrimental when consecutive decisions fall into distinct modes. Our method more effectively addresses cross-chunk dependency through dedicated behavioral search and is not mutually exclusive with the previous methods. We will demonstrate in the next section that combining our method with moving average can further improve closed-loop action chunking.
@@ 5 Experiments
    How does the proposed method affect the closed-loop operation of a policy with action chunking?
@@
    How does the proposed method scale with large batch sizes and complement existing methods?
To this end, we will first validate our theoretical analysis through one-dimensional diagnostic simulations. We will then evaluate BID on seven tasks in three simulation benchmarks, including Push-T [4], RoboMimic [56], and Franka Kitchen [57]. We will subsequently examine the generality and scalability of our method under various base policies and sample sizes. We will finally demonstrate the effectiveness of BID in two challenging real-world tasks involving dynamic objects.
5.1 One-Dimensional Diagnostic Simulations
Setup. We start with a diagnostic experiment in a one-dimensional state space $\{s_{0},s_{1},\cdots,s_{10}\}$, where $s_{0}$ is the starting state and $s_{10}$ is the goal state. The demonstrator plans to move forward by one step in each state, except in $s_{5}$ where it pauses unless the last five states visited were $s_{5}$. Each forward move has a success probability of $1-\delta$, where $\delta$ denotes the level of stochastic noise in the environment (as described in Sec. 3.1). Given these demonstrations, we train a collection of policies with different action horizons $h\in\{1,2,3,5,7,10\}$. We investigate under what action horizon our learner can better imitate the distribution of idle actions taken by the expert over multiple rollouts.
Result. As shown in Fig. 5, when the environment is deterministic ($\delta=0$), larger action horizons capture the expert distribution better, consistent with Corollary 2. With an action horizon of 10, the learner achieves zero total variation distance with the expert distribution. Conversely, when the environment is highly stochastic $\delta=0.8$, an action horizon of 1 outperforms all other learners. With moderate noise $\delta=0.4$, there is no discernible monotonic pattern due to the tradeoff revealed in Proposition 1. Refer to Appendix A for more detailed results.
@@ 5.1 One-Dimensional Diagnostic Simulations
5.2 Effects of BID on Closed-Loop Action Chunking
Setup. We next examine the effect of our decoding algorithm on closed-loop action chunking in seven simulation tasks. Throughout our experiments, we use Diffusion Policy [4], a state-of-the-art algorithm for generative behavioral cloning, trained on human demonstrations as the base policy. We use the official configurations and checkpoints of the Diffusion Policy in our experiments and consider several competitive inference methods as points of comparison:
@@ 5.2 Effects of BID on Closed-Loop Action Chunking
  • Vanilla [4]: Execute the first action of a randomly sampled sequence in a closed-loop manner.
  • Lowvar [13]: Similar to Vanilla, but reduce variance in the initial noise for the diffusion process.
@@
  • Warmstart [14]: Similar to Lowvar, but warm-start the initial noise for the diffusion process from the previous decision.
@@
  • Exponential Moving Average (EMA) [3]: Smooth action chunking by averaging a new prediction $a$ with the previous one $\hat{a}$ for each overlapping step $a_{t}=\lambda a_{t}+(1-\lambda){\hat{a}}_{t}$. This method is also known as temporal ensembling. Here, $\lambda\in(0,1)$ is the decay rate of the previous prediction. By default, we set $\lambda=0.75$.
For BID, we use batch size $N=30$ and mode size $K=10$. For each method-environment pair, we report an average score over $100$ episodes. Please refer to Appendix B for implementation details.
Result. Our main empirical finding is that while existing inference methods offer some benefits for closed-loop operations, they lack robustness. As shown in Table 1, Lowvar and Warmstart yield clear improvements over the vanilla closed-loop operation in specific tasks, such as Transport and Franka Kitchen. However, their average performance gains are relatively mild, likely due to the difficulty in controlling the prediction variance caused by stochastic noise at each step of the diffusion process. EMA generally produces better results, yet the improvements vary significantly across different tasks and even degrade the performance in Tool Hang. The challenges of tuning EMA are further discussed in Appendix A. In comparison, BID consistently achieves substantial gains across all tasks, surpassing the vanilla baseline by over $26\%$ in relative improvements.
Method Push T Lift Can Square Transport ToolHang Kitchen Average
Vanilla [4] 0.78 0.62 0.89 0.74 0.43 0.50 0.22 0.60
Lowvar [13] 0.79 0.54 0.91 0.75 0.52 0.52 0.25 0.61
Warmstart [14] 0.79 0.53 0.92 0.78 0.47 0.51 0.27 0.61
EMA [3] 0.83 0.77 0.92 0.75 0.59 0.46 0.51 0.69
BID (ours) 0.85 0.91 0.96 0.79 0.62 0.56 0.60 0.76
Table 1: Comparison of different methods for the closed-loop operation of diffusion policies. Evaluations are based on the mean score over $100$ episodes in the Push-T, RoboMimic, and 4-Object Franka Kitchen tasks.
5.3 Scalability and Compatibility of BID
@@
Setup. We further assess two key properties of BID: scalability with large batch sizes and compatibility with existing inference methods. For scalability, we experiment with batch sizes of $\{1,5,15,30\}$. For compatibility, we apply BID with a batch size of 15 to two competitive baselines, Warmstart and EMA. These experiments are conducted in the Push-T task.
@@
Result. As shown in Fig. 6, our method clearly benefits from the large batch size and is not yet saturated with the default batch size used in our experiment in Sec. 5.2. Moreover, the benefits from sampling with BID are complementary to that of existing inference methods. These two properties highlight the strong potential of the proposed method in practice.
Figure 6: BID benefits from more samples (left) and complements existing inference methods (right). All methods are evaluated for Diffusion Policy on the Pust-T task.
@@
5.4 Generality and Overhead of BID
Setup. To examine the generality and overhead of our method, we next extend our experiment to VQ-BET [6], a state-of-the-art transformer-based policy. Specifically, we use the public checkpoint on the Push-T task provided by LeRobot [58] as the base policy. We use a checkpoint early-terminated at 100 epochs as the weak policy in forward contrast. The computational time was measured on a desktop equipped with an NVIDIA A5000 GPU.
Result. Table 3 summarizes the results of the baseline and our method with a batch size of 16 samples. We observe that the vanilla random sampling performs significantly worse than BID in both closed and open-loop operations. Notably, the vanilla open-loop approach exhibits a rapid performance decline as the environment becomes increasingly stochastic. Even in closed-loop operations, the vanilla baseline still experiences a significant performance drop. In comparison, the closed-loop BID demonstrates much higher robustness to stochastic noise.
The experiments on VQ-BET also confirm the absence of a universally optimal action chunk size. Shorter action horizons tend to be more effective in noisy environments, while longer horizons excel in cleaner settings. This variability aligns with our theoretical analysis in Sec. 3.1.
Table 3 details the computational overhead associated with BID at varying batch sizes. The result shows that the performance gains of our method come with a 2-3x increase in computational overhead. We expect that this overhead will be less of a constraint with higher-end GPUs.
5.5 Real-world Experiments
Beyond the simulation experiments described above, we further evaluate the proposed BID through two real-world experiments.
5.5.1 Dynamic Placing

### v2

@@ 4 Method: Bidirectional Decoding
As analyzed above, action chunking improves long-term consistency but sacrifices short-term reactivity. In this section, we propose to address this tradeoff by bridging action chunking with closed-loop operations. We will first outline the general framework in Sec. 4.1 and then describe two specific criteria in Sec. 4.2.
@@ 4.1 Test-Time Search
Recall that for a policy with prediction horizon $l$, an action chunk $a:=\{a_{t}^{(t)},a_{t+1}^{(t)},\cdots,a_{t+l}^{(t)}\}$ sampled at time $t$ is expected to follow a consistent strategy over the subsequent $l$ time steps. However, executing the action chunk in an open-loop manner leaves the policy vulnerable to unexpected state changes. Alternatively, one can maximize reactivity by resampling an action chunk and executing only the first immediate action at every time step. Yet, this simple closed-loop approach destroys the consistency preserved within each chunk, potentially leading to oscillations between different strategies. How can we combine the benefits of both approaches?
Our key idea is to bridge action chunking and closed-loop resampling by making use of additional computes at test time. In particular, we seek to restore temporal consistency in closed-loop operations by scaling up the number of candidate samples. Intuitively, while the probability of any single pair of samples sharing the same latent strategy is low, the likelihood of finding a consistent pair increases with the number of samples. We thus frame the problem of closed-loop action chunking as searching for the optimal action among a batch of samples drawn at each time step,
@@
where $\mathcal{A}$ is the set of sampled action chunks, $\mathcal{L}_{B}$ and $\mathcal{L}_{F}$ are two criteria approximating the optimality with respect to the backward decision and forward plan, which we will describe next.
@@ 4.2 Bidirectional Criteria
Backward coherence. To preserve temporal dependencies in closed-loop operations, a sequence of actions should (i) commit to a consistent latent strategy over time, and (ii) react smoothly to unexpected changes. These desired properties motivate us to keep the action chunk selected at the previous time $\hat{a}:=\{{a}^{(t-1)}_{t-1},\cdots,{a}^{(t-1)}_{t+l-1}\}$ as a prior, and minimize the weighted Euclidean distance between the new action chunk and the prior across $l-1$ overlapping steps:
@@
Here, $\rho$ is a decay hyperparameter to account for growing uncertainty over time. This backward objective encourages similar latent strategies between consecutive steps while allowing for gradual adaptation to unforeseen transition dynamics.
Nevertheless, the backward criterion alone presents a potential caveat: the prior chunk could be suboptimal due to the lack of information at the previous time step (e.g., unexpected object motions). In such cases, selecting the next action chunk based solely on the prior may perpetuate suboptimality. Ideally, the sequential decision-making process should effectively correct suboptimal plans based on the latest observations. We next address this need through another forward criterion.
Figure 4: Illustration of bidirectional decoding.
@@
Forward contrast. Our design of the forward criterion is motivated by the need to identify the most optimal plan from a set of candidates. Within the same latent strategy, suboptimal samples may arise from (i) low likelihood under the learned model and (ii) divergence between the learned model and expert policy. To address this, we draw inspirations from LLM decoding techniques (Wang et al., 2022; Li et al., 2023) and introduce a forward contrast criterion. Specifically, we compare each candidate sample with two sets of reference samples: one positive set from a stronger policy and a negative set from a weaker one. The stronger policy is obtained from a well-trained checkpoint, whereas the weaker policy is taken from an early underfitting checkpoint and is expected to be further from the expert policy. Our forward objective is thus framed as minimizing the average distance between a candidate plan and the positive samples while maximizing its average distance from the negative ones,
Fig. 4 illustrates the combined effects of the backward coherence and forward contrast criteria on sample selection. Since samples in $\mathcal{A}^{+}$ and $\mathcal{A}^{-}$ are not all subject to the same strategy, we trim each set by removing samples that deviate significantly from the previous decision. This is achieved by summing over the $K$ smallest distance values for in the positive and negative sets in Eq. 7. The full process of our decoding method is outlined in Algorithm 1. Since all steps in BID can be computed in parallel, the overall computational overhead remains modest on modern GPU devices.
@@ 4.3 Discussions
Interpretation of our method. Our method makes no changes to the learned policy; instead, it intervenes in the model distribution through sample selection. As illustrated in Fig. 10, randomly sampled sequences may be misaligned with both the previous decisions and the target demonstrations. Given a set of candidates, the backward step first identifies the behavioral mode from the past decision stored in memory; the forward step then removes the samples with low likelihood under the target distribution using prior knowledge of positive and negative samples. By comparing samples across time steps and model checkpoints, our method bridges the gap between the proposal and target distributions during inference.
Relation to recent methods. Our method builds upon the receding horizon (Chi et al., 2023) and temporal ensembling (Zhao et al., 2023) used in previous works, but with crucial distinctions. Receding horizon seeks a compromise between long-term consistency and short-term reactivity by using a moderate action horizon (e.g., half of the prediction horizon), which is inevitably sup-optimal when both factors are prominent. Temporal ensembling strengthens dependency across chunks by averaging multiple decisions over time; however, weighted-averaging operations can be detrimental when consecutive decisions fall into distinct modes. Our method more effectively addresses cross-chunk dependency through dedicated behavioral search and is not mutually exclusive to previous methods. We will demonstrate in the next section that combining our method with the moving average can further improve closed-loop action chunking.
@@ 5 Experiments
    Can our decoding method improve closed-loop operations of policies built with action chunking?
@@
    Does our decoding method perform well across different policies, tasks, and environments?
  4. 4.
    Is our decoding method scalable to larger sample sizes and compatible with existing methods?
@@
To this end, we will first validate our theoretical analysis through one-dimensional diagnostic simulations. We will then evaluate BID on seven tasks across three simulation benchmarks, including Push-T (Chi et al., 2023), RoboMimic (Mandlekar et al., 2022), and Franka Kitchen (Gupta et al., 2020). We will subsequently examine the generality and scalability of our method under various base policies and sample sizes. We will finally assess the effectiveness of BID in two challenging real-world tasks that require interactions with dynamic objects.
5.1 One-Dimensional Diagnostic Experiments
Setup. We start with a diagnostic experiment in a one-dimensional state space $\{s_{0},s_{1},\cdots,s_{10}\}$, where $s_{0}$ is the starting state and $s_{10}$ is the goal state. The demonstrator plans to move forward by one step in each state, except in $s_{5}$ where it pauses unless the last five states visited were $s_{5}$. Each forward move has a probability of $1-\delta$, where $\delta$ denotes the level of stochastic noise in the environment (as described in Sec. 3.2). Given these demonstrations, we train a collection of policies with different action horizons $h\in\{1,2,3,5,7,10\}$. We investigate under what action horizon our learner can better imitate the distribution of idle actions taken by the expert over multiple rollouts.
Result. As shown in Fig. 5, when the environment is deterministic ($\delta=0$), larger action horizons capture the expert distribution better, consistent with Corollary 2. With an action horizon of 10, the learner achieves zero total variation distance with the expert distribution. Conversely, when the environment is highly stochastic $\delta=0.8$, an action horizon of 1 outperforms all other learners, corroborating with Corollary 3. With moderate noise $\delta=0.4$, there is no discernible monotonic pattern due to the tradeoff revealed in Proposition 1. Refer to Appendix A for more detailed results.
5.2 Simulation Experiments with Stochastic Noise
Next, we evaluate our decoding algorithm on seven simulation tasks of robot manipulation. We will first compare BID with existing inference methods in closed-loop operations. We will then assess the effectiveness of our method under different conditions, including policy classes, sample sizes, and levels of stochastic noise.
5.2.1 Comparison with Existing Inference Methods
Setup. In each manipulation task, we use Diffusion Policy (Chi et al., 2023) trained on human demonstrations as the base policy. We evaluate BID with a batch size of $N=16$ and a mode size of $K=3$. We consider three existing inference methods as baselines:
@@ 5.2.1 Comparison with Existing Inference Methods
  • Vanilla (Chi et al., 2023): Execute the first action of a sampled chunk in a closed-loop manner.
@@
  • Warmstart (Janner et al., 2022): Similar to Vanilla, but warm-start the initial noise for the diffusion process from the previous decision.
@@
  • Exponential Moving Average (EMA) (Zhao et al., 2023): Smooth action chunking by averaging a new prediction $a$ with the previous one $\hat{a}$ for each overlapping step $a_{t}=\lambda a_{t}+(1-\lambda){\hat{a}}_{t}$. This method is also known as temporal ensembling. By default, we set $\lambda=0.5$.
We evaluate each method for $100$ episodes and average the results across three random seeds. Please refer to Appendix B for implementation details.
Figure 6: Comparison of different inference methods for closed-loop operations of diffusion policies. Each method is evaluated for $100$ episodes on seven manipulation tasks in simulation benchmarks. Results are averaged across three seeds. BID significantly outperforms existing inference methods in most tasks.
Figure 7: BID benefits from large sample sizes (left) and complements existing inference methods (right). Each method is evaluated on seven simulation tasks across three seeds. Relative performance gain is measured with respect to the vanilla baseline. When combined with EMA, BID results in 46% relative improvements.
Result. Our main observation is that while existing inference methods offer some benefits for closed-loop operations, they lack robustness. As shown in Fig. 6, Warmstart yields mild performance gains on average, but degrades performance on 3 out of 7 tasks. Similarly, EMA leads to competitive results on several tasks, yet exhibits performance drops in 2 tasks. We conjecture that this robustness issue stems from independent sampling across chunks; when successive chunks follow distinct latent strategies, averaging them may not yield a plausible strategy, as further discussed in Sec. A.4. In comparison, BID offers substantial gains across all tasks. Notably, BID provides $32\%$ relative improvements over the vanilla baseline, significantly outperforming EMA on Pust-T, Lift, Square, and Tool Hang, while achieving competitive performance on the other tasks.
5.2.2 Scalability and Compatibility of BID
@@
Setup. We further assess two key properties of BID: scalability with growing batch sizes and compatibility with existing inference methods. For scalability, we experiment with batch sizes of $\{1,4,8,12,16\}$. For compatibility, we apply BID on top of two other baselines, Warmstart and EMA. The results are averaged across seven simulation tasks and three random seeds.
@@
Result. As shown in Fig. 7, our method benefits from the large batch size, with performance gains not yet saturated at the default batch size used in Sec. 5.2.1. Moreover, the strength from BID is complementary to that of existing inference methods. Notably, combining BID with EMA further boosts the relative performance gain from 32% to 46%. These two properties highlight the potential of our method in practice.
5.2.3 Generality and Efficiency of BID
Setup. We next extend our experiment to VQ-BET (Lee et al., 2024), another state-of-the-art robot policy built with autoregressive transformers. We use the public checkpoint on the Push-T task provided by LeRobot (Cadene et al., 2024) as the base policy. We use a checkpoint terminated at 100 epochs as the weak policy in forward contrast. To simulate stochastic conditions, we add temporally correlated Gaussian noise to the executed action at each step, scaled by the action magnitude. We measure the computational time on a desktop equipped with an NVIDIA A5000 GPU.
Result. Table 2 summarizes the results of the baseline and our method. The vanilla random sampling performs significantly worse than BID in both closed and open-loop operations. Notably, the vanilla open-loop approach exhibits a rapid performance decline as the environment becomes increasingly stochastic. Even in closed-loop operations, the vanilla baseline still experiences a significant performance drop. In comparison, the closed-loop BID demonstrates much higher robustness to stochastic noise. Table 2 details the computational overhead associated with BID at varying batch sizes. The result shows that the performance gains of our method come with a doubled computational overhead. We expect that this overhead will be less of a constraint with higher-end GPUs.
@@ 5.2.3 Generality and Efficiency of BID
5.3 Real-world Experiments with Dynamic Objects
We finally evaluate BID through two real-world experiments that require rapid reactions to dynamically moving objects.
Setup. We consider two pick-and-place tasks, where the target object undergoes unexpected movement during evaluation. In the dynamic placing task, a Franka Panda robot is required to deliver an object into a moving cup held by a human subject. In the dynamic picking task, a UR5 robot is required to grasp a moving cup pulled by a string and place it onto a nearby static saucer. In both tasks, we evaluate the performance of BID applied to pre-trained diffusion policies. For further experimental details, please refer to Sec. B.2.
Result. Figs. 9 and 9 compares the results of different inference methods on the two real-world tasks. Vanilla random sampling struggles to handle the diverse demonstrations and dynamic movements, resulting in significantly lower success rates. In contrast, BID achieves high success rates in both static and dynamic conditions. Notably, in the dynamic picking task, BID achieves a 2x higher success rate than all other baselines, highlighting its potential for dynamic object interactions.
Other experiments.
Please refer to Appendix A for additional analyses and ablations.
6 Conclusion
Reproducibility Statement
To facilitate the reproducibility of our results, we release our code for the experiments with Diffusion Policy at https://github.com/YuejiangLIU/bid_diffusion and our code for the experiments with VQ-BET at https://github.com/Jubayer-Hamid/bid_lerobot. Additionally, we provide our detailed experimental setups in Appendix B, and complete proofs of our theoretical analysis in Appendix C.
Acknowledgments
Appendix A Additional Experiments
A.1 One-dimensional Simulations
Action Horizon
Noise Level 1 3 5 7 10
0.0 4.21 1.75 1.55 1.28 0.00
0.4 0.55 0.30 0.95 0.53 0.93
0.8 0.04 0.98 1.23 1.26 1.44
A.2 Action Horizon vs. Context Length
Setup. Our work builds on the premise that the action horizon is longer than the context length, as commonly designed for recent policies. While BID mitigates the inherent limitations of this design choice through test-time decoding, an important question remains: could extending the context length yield stronger policies? To understand this, we trained diffusion policies with varying combinations of action horizons and context lengths on the Push-T task. Specifically, we use a short context length ($c=2$) and a short action horizon ($h=2$) as our baseline, and incrementally increase these parameters to larger values ${6,10,14}$ to assess their impact.
Result. Fig. 11 compares the performance of the policy learned with different $\Delta h=h-c$. As expected, the policy with both a short action horizon and a short context length does not perform well, due to its limited capability to model long-range temporal dependencies. Interestingly, extending the context length initially boosts performance ($\Delta h=-4$), but this trend reverses as the context length becomes too long ($\Delta h\leq-8$), likely due to overfitting to an increased number of spurious features. In contrast, expanding the action horizon results in more robust performance improvements, validating its pivotal role in imitation learning from human demonstrations.
A.3 Ablation Study of Forward Contrast
Setup.
To understand the effect of forward contrast (Equation 7), we evaluate the full version of our method against three reduced variants in open-loop operations: Vanilla (without forward contrast), Positive (without negative samples), and Negative (without positive samples). Similar to Sec. 5.2.2, our ablation study is conducted in seven simulation tasks, each with three random seeds.
Result.
Fig. 13 summarizes the result of this ablation study. Notably, both positive and negative samples are essential for effective sample selection, and omitting either leads to significant performance declines. Without negative samples, our decoding method reduces to an approximate maximum a posteriori estimation, which can result in suboptimal decisions due to modeling errors. Conversely, without positive samples, the sampling process may be biased towards rare instances. This result highlights the importance of both components and suggests the potential for extending this paradigm in future work.
Figure 12: Effect of positive and negative samples on forward contrast. Performance of ablated variants of forward contrast is evaluated across seven simulation tasks. The absence of either positive or negative samples prevents achieving the full performance gains observed with the contrast objective.
Figure 13: Effect of the decay rate for the exponential moving average. In each task, we measure the relative performance among different decay rates. The optimal decay rate varies by task, leading to a practical challenge of identifying a universal temporal ensembling strategy (Zhao et al., 2023).
A.4 Challenges for Temporal Ensebmling
EMA exhibits competitive performance in Fig. 6. However, tuning its decay rate can be difficult in practice. Fig. 13 shows the sensitivity of EMA to the decay rate across three different tasks, where the optimal choices differ significantly. We conjecture that this high sensitivity stems from the variability in the latent strategies between consecutive predictions. When consecutive predictions follow similar strategies, a lower decay rate (i.e., stronger moving average) can enhance smoothness and improve performance. Conversely, when consecutive predictions diverge in their underlying strategies, averaging them can introduce adverse effects. Our method promotes coherence in latent strategies and thus effectively complements temporal ensembling, as evidenced in Fig. 7.
Appendix B Additional Details
B.1 Simulation Experiment Details
B.1.1 Environment Details
Push-T: We adopt the Push-T environment introduced in (Chi et al., 2023), where the goal is to push a T-shaped block on a table to a target position. The action space is two-dimensional end-effector velocity control. The training dataset contains 206 demonstrations collected by humans.
Robomimic: We use five tasks in the Robomimic suite (Mandlekar et al., 2022), namely Lift, Can, Square, Transport, and Tool Hang. The training dataset for each task contains 300 episodes collected from multi-human (MH) demonstrations.
Franka Kitchen: We use the Franka Kitchen environment from (Gupta et al., 2020), featuring a Franka Panda arm with a seven-dimensional action space and 566 human-collected demonstrations. The learned policy is evaluated on test cases involving four or more objects (p4), a challenging yet practical task for robotic manipulation in household contexts.
B.1.2 Implementation Details.
Our implementation of BID for Diffusion Policy is built upon the official code of Chi et al. (2023), with modifications made solely to the inference process. The policy takes in state inputs and predicts a chunk of 16 actions as outputs. For each simulation task, we train the model for 100-1000 epochs to reach near-optimal performance. We evaluate it in closed-loop operations, i.e., action horizon is set to 1. For forward contrast, we train the weak policy for 10-100 epochs, resulting in a suboptimal policy for each task. The core hyperparameters are summarized in Table 4.
Our implementation of BID for VQ-BeT (Lee et al., 2024) is built upon the code of LeRobot (Cadene et al., 2024). We use the best public checkpoint as the strong policy and a checkpoint trained for 100k iterations as the weak policy. Since BID requires sample diversity, we set the temperature as 0.5 for our methods, while vanilla sampling uses the default temperature of 0.1.
name value
batch size $N$ 16
mode size $K$ 3
prediction horizon $l$ 16
temporal coherence decay $\rho$ 0.5
moving average decay $\lambda$ 0.5
Table 4: Default hyper-parameters in our experiments.
B.2 Real-world Experiment Details
B.2.1 Dynamic Placing